As a coin on a turntable spins in a circle with speed 8.0 m/s it experiences a centripetal acceleration of 2.0 m/s^2. What is the radius and period?
Answer by mananth(16949)

centripetal acceleration = v^2/r r= radius
v = 8 m/s
a= 2m/s^2
r=v^2/a
r= 64/2
r= 32 m